• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name M Friend

    Log entry time 00:12:10 on October27,2009

    Entry number 296846

    keyword=Compton run summary

    Run 20680: 5uA coincidence run with edet at -3.0mm. Prescales 4%,100%, 100%. 2 plane coincidences. Doesn't seem to be working.

    Run 20681: 5uA, electrons only, FADC trigger messed up, edet at -3.0mm, 100% electrons, 2 plane coincidences.

    Run 20682: 10uA, coinc run, FADC normal trigger, edet at -3.0mm, 100% everything, 2 plane coincidences. Doesn't work.

    Run 20683: Reduced FIFO from 300 to 100 for edet, 200 to 100 for photon. Another 10uA coincidence run with prescales 100%. Still doesn't work.

    Run 20684: photons+electrons+coincidence. Still at 10uA, still with fifo changes, still 100% prescales. There are electrons now.

    Run 20688: electrons only. 1uA.

    Run 20689: 10uA, electron only, FADC trig on edet, reloaded firmware. Crashed in the middle.

    Run 20694: electron only, FADC trig on edet.

    Run 20695: coincidence run. 2.5uA (doesn't work) then 5uA where it does work.

    Maybe we didn't wait long enough for the other runs.

    More coincidence runs at different currents (100% for all prescales):

    20696 (5uA)
    20699 (10uA)